OpenStudy (anonymous):

Solve by substitution -3x-3y=3 y=-5x-17

OpenStudy (anonymous):

-3x - 3(-5x-17) = 3 -3x + 15x +51 = 3 12x = 3-51 12x = -48 x = -4 y = -5(-4) -17 y = 20 -17 y = 3

OpenStudy (anonymous):

I wish I could give you more medals:)

OpenStudy (anonymous):

that is sweet :) believe me iam so hooked onto this site and having sso much fun.